OpenStudy (anonymous):

PLEASE HELP write an equation in standard form of the line that has the given properties. slope 1/2 y-intercept 2 slope 1 point (2,-3) please show work ill give you best answer medal

OpenStudy (anonymous):

should be y=1/2x+2 and y=1x-5

OpenStudy (anonymous):

i used y=mx+b btw

OpenStudy (anonymous):

wait I don't understand it ask me to do it in standard form how did you use point slope form ? @zach1751

OpenStudy (anonymous):

and where did you get -5 from ?

OpenStudy (anonymous):

im sorry im getting tired i just totally like read over that part without giving it a thought

OpenStudy (anonymous):

it didnt make sense to me either but -5 was the only way to get it to pass through the (2,-3)

OpenStudy (anonymous):

okay so haw do I write an equation in point slope from ?

OpenStudy (anonymous):

do you know how to turn point slope into standard?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok you use the template y=mx+b because thats all it really is is a template. ok, now m represents your slope so you want a slope of one you put 1 for m. dont worry about x that stays the same. b is where the line intersects the y-axis. ok?

OpenStudy (anonymous):

does this make sense?

OpenStudy (anonymous):

i understand that but standard form is AX+BY=C

OpenStudy (anonymous):

yes if you take the equation y=1x-5 you should just be able to move the 1x over so... -1x+y=-5 could you not?

OpenStudy (anonymous):

Ohhhhhhh lol thanks you and how bout if i have points (3,4) and (-1,0) ?

OpenStudy (anonymous):

i think the -5 should be positive from what i look at online you should divide EVERYTHING by -1 which reverses all of their signs therefore it is x-y=5

OpenStudy (anonymous):

for your other question it should be x-y=-1

OpenStudy (anonymous):

why divide everything by -1?

OpenStudy (anonymous):

its just one of things that you do lol

OpenStudy (anonymous):

lol okay thank you so much i just need help with one more problem can you help me please

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay thank you. write an equation of the line in slope intercept form that passes through (-2,3) and is parallel to the line y=2x+1

OpenStudy (anonymous):

im sorry i did not notify me that you responded

OpenStudy (anonymous):

do you have any ideas?

OpenStudy (anonymous):

ok i have to go the answer is y=2x+7 i wish i had time to explain but if you still dont understand by tommorow just private message me.